You are here: Foswiki>AdaptiveOptics Web>Logs>Lucifer217FebFeb_2014 (20 Feb 2014, JuanGuerra)Edit Attach

UT 20140217UT RR+Lucifer2 DX

Observer:
Telescope Operator:
SW Support:k
AO Support: r
Telescope Support:
Instrument Support:

Work

Installation of the RR to test the new firmware on the SwitchBCU. The FLAO system will be used to close the loop and to check the new firmware before LBTI checks. Marco Xompero is available on Skype after 11am(Tucson time)

Summary

  • Focal station RFBG.
  • Lucifer 2 is install.
  • RR installed about 8 am. Telescope at zenith with HBS on and stop pin IN. Floor light on.
  • Argos dichroic support installed infront of the AGW.
  • Power on WFS with out any problem.
  • Acquiring the return.
  • Adsec DX ok and Flat.
  • Authorize the RR focal station: RETROREF.
  • Spare SwitchBCU memory flash and new firmware loaded and tested with non success result.
  • Handover 4pm
  • Retro-reflector off

WFS

The WFS was power on without any problem. The system had an issue with the slope computer that lost communication 4 times in one hour. After that the system runs without any further problems. It seems that the electronics requires about one hour to be in operating temperature. Warmup period 1 hour. The internal source was acquired and it was necessary to move the stages to see the internal source. the positions for the BAyside stage are: x=-77.5mm,y=-50mm,z=40 (nominal)

Acquisition of the return.

The telescope is authorized with RETROREF configuration. This is the configuration for the RFBG to work with the RR.

The absolute values are:

Hexapod:
x(mm) y(mm) z(mm) RX(") Ry(") Rz(") Note
1.89 2.22 5.05 -384 265.98 0

M3: global position:
Tip Tilt Piston SelAngle
172 349 0 0
Absolute position:
Tip Tilt Piston SelAngle
-386 679 0 -25.83

With this values on, we couldn't find the spot and we moved back to the October RR positions for M3.

The spot looks vignetting on the left side, moving the bayside state the vignetting didn't move and after moving 20" Tip on the tertiary the spot ca be seeing without the vignetting To compensate an astigmatism we applied -2um on Z6 from the AdsecEng Gui.

A spiral search around of the position didn't find the second reflection from the dichoric, ( is there a second reflection?)

Issues:

  • The slope computer has stopped four time in the first hour that the unit is power on.

20141702 15:09:57.6867 Alert: Error: LOST COMMUNICATION WITH SLOPE COMPUTER
20141702 15:11:53.8887 Alert: Error: LOST COMMUNICATION WITH SLOPE COMPUTER
20141702 15:34:01.8950 Alert: Error: LOST COMMUNICATION WITH SLOPE COMPUTER
20141702 15:53:26.5479 Alert: Error: LOST COMMUNICATION WITH SLOPE COMPUTER

The slope computer has not showed more problems with lost communications. It seems that it needs about one hour. Warm up !!

  • Internal source is not visible. The configuration for bayside stage and cube has changed. The script

The spot is found at:

Bayside state x=-74.5 y=-50 z=40 (nominal)
Cube rot 181.65 (penina)
Cube stage -4.4 (nominal.. IN)

  • Pupil vignetting in one side by a long bar crosing the detector.

Looking the pupil we can see that there is a vignetting on the left side of the CCD47. See image. Moving the Teriary by 20" in Tip and moving the stage in X, the vignetting in not present. See image. This could be an issue of the ARGOS dichroics structure infornt of the AGW?

HEXAPOD and M3 POSITIONS:

The telesacope is authorize to RETROREFF (RETRO-REFLECTOR RIGHT FRONT BENT GREGORIAN)

  • HEXAPOD GLOBAL OFFSETS

x(mm) y(mm) z(mm) RX(") Ry(") Rz(") Note
1.89|2.22|5.1|-385|295||

  • M3 GLOBAL OFFSET:

Tip(") tilt(") piston Sel angle(deg) note
172 349 0 0

bayside stage position:

x y z
-76.3 -52.3 40

Swap Switch BCU

The spare SWBCU was flashed with the new firmware from Mario and it has been tested on the telescope. The result showed that the system can not open the fastlink between BU and the fastlink check fails. A file is attached with some of the errors: * log_switchbcu: LOG_BCU


adsec.R |INF| 361|2014-02-17 23:52:36.932946| MAIN > FastLink connection testing... adsec.R |INF| 362|2014-02-17 23:52:36.954044| MAIN > Fastlink alignment procedure initialized. adsec.R |WAR| 363|2014-02-17 23:52:37.090446| MAIN > [RequestFaultException] The request has failed adsec.R |WAR| 364|2014-02-17 23:52:37.091154| MAIN > [RequestFaultException] The request has failed adsec.R |INF| 365|2014-02-17 23:52:37.222141| MAIN > Fastlink test failed but unable to find the DSP which failed.
mirrorctrl.R |ERR| 567|2014-02-17 23:52:37.090056| MAIN > ****** FAILED PACKET FROM BCU 2 - CMD TIMEOUT CA ***** mirrorctrl.R |ERR| 568|2014-02-17 23:52:37.090212| MAIN > ****** FAILED PACKET FROM BCU 2 - CMD TIMEOUT CA ***** mirrorctrl.R |ERR| 569|2014-02-17 23:52:37.090251| BCUCOMMANDHANDLE > BcuCommandHandler::handleBcuReply: REQUEST COMPLETED but FAILED frown, sad smile mirrorctrl.R |ERR| 570|2014-02-17 23:52:37.090947| MAIN > ****** FAILED PACKET FROM BCU 2 - CMD TIMEOUT CA *****


CCD39 RON test


SUMMARY ------------------------------------
Bin: 1 - Framerate 990 Hz (2500 kpix/sec) - RON (e-): 10.88 (estimated)
Bin: 2 - Framerate 990 Hz ( 800 kpix/sec) - RON (e-): 6.03 (estimated)
Bin: 2 - Framerate 625 Hz ( 380 kpix/sec) - RON (e-): 4.35 (estimated)
Bin: 2 - Framerate 400 Hz ( 380 kpix/sec) - RON (e-): 4.30 (estimated)
Bin: 3 - Framerate 200 Hz ( 380 kpix/sec) - RON (e-): 4.52 (estimated)
Bin: 4 - Framerate 300 Hz ( 335 kpix/sec) - RON (e-): 4.37 (estimated)
Bin: 4 - Framerate 200 Hz ( 335 kpix/sec) - RON (e-): 4.50 (estimated)
Bin: 4 - Framerate 105 Hz ( 335 kpix/sec) - RON (e-): 4.53 (estimated)


SUMMARY ------------------------------------
Bin: 1 - Framerate 990 Hz (2500 kpix/sec) - RON (e-): 10.67 (estimated)
Bin: 2 - Framerate 990 Hz ( 800 kpix/sec) - RON (e-): 5.97 (estimated)
Bin: 2 - Framerate 625 Hz ( 380 kpix/sec) - RON (e-): 4.21 (estimated)
Bin: 2 - Framerate 400 Hz ( 380 kpix/sec) - RON (e-): 4.34 (estimated)
Bin: 3 - Framerate 200 Hz ( 380 kpix/sec) - RON (e-): 4.30 (estimated)
Bin: 4 - Framerate 300 Hz ( 335 kpix/sec) - RON (e-): 4.45 (estimated)
Bin: 4 - Framerate 200 Hz ( 335 kpix/sec) - RON (e-): 4.40 (estimated)
Bin: 4 - Framerate 105 Hz ( 335 kpix/sec) - RON (e-): 4.50 (estimated)

The fastlink is checked using the close loop at gain zero. The loop close without problem and we don't see any CA timeout in the BCUinterface logs:

Starting: 2014-02-17 16:37:56.188833| MAIN > arbFunc doCloseLoop done. End: |2014-02-17 16:47:30.193960| MAIN > arbFunc doStopLoop done. Start: WfsArbitrator_22037|DEB| 149|2014-02-17 16:50:06.562260| MAIN > arbFunc doCloseLoop done. End:1|2014-02-17 16:56:45.097191| MAIN > arbFunc doStopLoop done.

-- JuanGuerra - 17 Feb 2014
Topic revision: r5 - 20 Feb 2014, JuanGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback